# Dual-Objective Reinforcement Learning with Novel Hamilton-Jacobi-Bellman Formulations

### v1

@@ VIII Conclusion
To improve our algorithm, we incorporate the Generalized Advantage Estimate corresponding to these Bellman equations in the updates of the Actors. As outlined in Section A of [21], the GAE may be defined with a reduction function corresponding to the appropriate Bellman function which will be applied over a trajectory roll-out. We generalize the Reach GAE definition given in [21] to propose a Reach-Avoid GAE (the Avoid GAE is simply the flip of the Reach GAE) as all will be used in DO-HJ-PPO algorithm for either RAA or RR problems. Consider a reduction function $\phi_{RA}^{(n)}:\mathbb{R}^{n}\to\mathbb{R}$, defined by
@@
which may be approximated over any finite trajectory sample. See [21] for further details.
-F2 Modifications from standard PPO
To address the RAA and RR problems, DO-HJ-PPO introduces several key modifications to the standard PPO framework [23]:
@@
Rather than learning the decomposed objectives separately from the composed objective, DO-HJ-PPO optimizes all objectives simultaneously. This design choice is motivated by two primary factors: (i) simplicity and minor computational speed-up, and (ii) coupling between the decomposed and composed objectives during learning.
@@
While it is possible to estimate the decomposed objectives independently—i.e., prior to solving the composed task—this approach might lead to inaccurate or irrelevant value estimates in on-policy settings. For example, in the RAA problem, the decomposed objective may prioritize avoiding penalties, while the composed task requires reaching a reward region without incurring penalties. In such a case, a decomposed policy trained in isolation might converge to an optimal strategy within a reward-irrelevant region, misaligned with the overall task. Empirically, we observe that omitting coupled resets causes DO-HJ-PPO to perform no better than standard baselines such as CPPO, whereas their inclusion significantly improves performance.
@@
-G DDQN Demonstration
As described in the paper, we demonstrate the novel RAA and RR problems in a 2D $Q$-learning problem where the value function may be observed easily. We juxtapose these solitons with those of the previously studied RA and R problems which consider more simple objectives. To solve all values, we employ the standard Double-Deep $Q$ learning approach (DDQN) [25] with only the special Bellman updates.
-G1 Grid-World Environment
The environment is taken from [19] and consists of two dimensions, $s=(x,y)$, and three actions, $a\in\{\text{left},\text{straight},\text{right}\}$, which allow the agent to maneuver through the space. The deterministic dynamics of the environment are defined by constant upward flow such that,
@@
-G2 DDQN Details
As per our theoretical results in Theorems 1 and 2, we may now perform DDQN to solve the RAA and RR problems with solely the previously studied Bellman updates for the RA [19] and R problems [18]. We compare these solutions with those corresponding to the RA and R problems without the special RAA and RR targets, and hence solve the previously posed problems. For all experiments, we employ the same adapted algorithm as in [19], with no modification of the hyper-parameters given in Table I.
TABLE I: Hyperparameters for DDQN Grid World
@@
-H Baselines
In both RAA and RR problems, we employ Constrained PPO (CPPO) [2] as the major baseline as it can handle secondary objectives which are reformulated as constraints. The algorithm was not designed to minimize its constraints necessarily but may do so in attempting to satisfy them. As a novel direction in RL, few algorithms have been designed to optimize max/min accumulated costs and thus CPPO serves as the best proxy. Below we also include a naively decomposed STL algorithm to offer some insight into direct approaches to optimizing the max/min accumulated reward.
-H1 CPPO Baselines
@@
-I Details of RAA & RR Experiments: Hopper
The Hopper environment is taken from Gym [26] and [21]. In both RAA and RR problems, we define rewards and penalties based on the position of the Hopper head, which we denote as $(x,y)$ in this section.
In the RAA task, the reward is defined as
$r(x,y)=\sqrt{\|x-2\|+|y-1.4\|}-0.1$ (51)
In the RR task, the first reward is defined again as
$r_{1}(x,y)=\sqrt{\|x-2\|+|y-1.4\|}-0.1$ (52)
$r_{2}(x,y)=\sqrt{\|x-0\|+|y-1.4\|}-0.1$ (53)
In all experiments, the Hopper is initialized in the default standing posture at a random $x\in[0,2]$ so as to learn a position-agnostic policy. The DO-HJ-PPO parameters used to train these problems can be found in Table II.
TABLE II: Hyperparameters for Hopper Learning
Hyperparameters for DO-HJ-PPO Values
@@
-J Details of RAA & RR Experiments: F16
The F16 environment is taken from [21], including a F16 fighter jet with a 26 dimensional observation. The jet is limited to a flight corridor with up to $2000$ relative position north ($x_{PN}$), $1200$ relative altitude ($x_{H}$), and ±$\pm 500$ relative position east ($x_{PE}$).
@@
In all experiments, the F16 is initialized with position $x_{PN}\in[250,750],x_{H}\in[300,900],x_{PE}\in[-250,250]$ and velocity in $v\in[200,450]$. Additionally, the roll, pitch, and yaw are initialized with ±$\pm\pi/16$ to simulate a variety of approaches to the flight corridor. Further details can be found in [21]. The DO-HJ-PPO parameters used to train these problems can be found in Table III.
TABLE III: Hyperparameters for F16 Learning
Hyperparameters for DO-HJ-PPO Values
@@
-K Broader Impacts
@@
-L Acknowledgments
We would like to thank Kai-Chieh Hsu and Oswin So for sharing their code bases which were fundamental to the success of the project.

### v2

@@ E.1 The special Bellman updates and the corresponding GAEs
To improve our algorithm, we incorporate the Generalized Advantage Estimate corresponding to these Bellman equations in the updates of the Actors. As outlined in Section A of So et al. [2024], the GAE may be defined with a reduction function corresponding to the appropriate Bellman function which will be applied over a trajectory roll-out. We generalize the Reach GAE definition given in So et al. [2024] to propose a Reach-Avoid GAE (the Avoid GAE is simply the flip of the Reach GAE) as all will be used in DOHJ-PPO algorithm for either RAA or RR problems. Consider a reduction function $\phi_{RA}^{(n)}:\mathbb{R}^{n}\to\mathbb{R}$, defined by
@@
which may be approximated over any finite trajectory sample. See So et al. [2024] for further details.
E.2 Modifications from standard PPO
To address the RAA and RR problems, DOHJ-PPO introduces several key modifications to the standard PPO framework Schulman et al. [2017]:
@@ E.2 Modifications from standard PPO
Rather than learning the decomposed objectives separately from the composed objective, DOHJ-PPO optimizes all objectives simultaneously. This design choice is motivated by two primary factors: (i) simplicity and minor computational speed-up, and (ii) coupling between the decomposed and composed objectives during learning.
@@
While it is possible to estimate the decomposed objectives independently—i.e., prior to solving the composed task—this approach might lead to inaccurate or irrelevant value estimates in on-policy settings. For example, in the RAA problem, the decomposed objective may prioritize avoiding penalties, while the composed task requires reaching a reward region without incurring penalties. In such a case, a decomposed policy trained in isolation might converge to an optimal strategy within a reward-irrelevant region, misaligned with the overall task. Empirically, we observe that omitting coupled resets causes DOHJ-PPO to perform no better than standard baselines such as CPPO, whereas their inclusion significantly improves performance.
@@
Appendix F DDQN Demonstration
As described in the paper, we demonstrate the novel RAA and RR problems in a 2D $Q$-learning problem where the value function may be observed easily. We juxtapose these solitons with those of the previously studied RA and R problems which consider more simple objectives. To solve all values, we employ the standard Double-Deep $Q$ learning approach (DDQN) Van Hasselt et al. [2016] with only the special Bellman updates.
F.1 Grid-World Environment
The environment is taken from Hsu et al. [2021] and consists of two dimensions, $s=(x,y)$, and three actions, $a\in\{\text{left},\text{straight},\text{right}\}$, which allow the agent to maneuver through the space. The deterministic dynamics of the environment are defined by constant upward flow such that,
@@ F.1 Grid-World Environment
F.2 DDQN Details
As per our theoretical results in Theorems 1 and 2, we may now perform DDQN to solve the RAA and RR problems with solely the previously studied Bellman updates for the RA Hsu et al. [2021] and R problems Fisac et al. [2019]. We compare these solutions with those corresponding to the RA and R problems without the special RAA and RR targets, and hence solve the previously posed problems. For all experiments, we employ the same adapted algorithm as in Hsu et al. [2021], with no modification of the hyper-parameters given in Table 1.
Table 1: Hyperparameters for DDQN Grid World
@@ F.2 DDQN Details
Appendix G Baselines
In both RAA and RR problems, we employ Constrained PPO (CPPO) Achiam et al. [2017a] as the major baseline as it can handle secondary objectives which are reformulated as constraints. The algorithm was not designed to minimize its constraints necessarily but may do so in attempting to satisfy them. As a novel direction in RL, few algorithms have been designed to optimize max/min accumulated costs and thus CPPO serves as the best proxy. Below we also include a naively decomposed STL algorithm to offer some insight into direct approaches to optimizing the max/min accumulated reward.
G.1 CPPO Baselines
@@ G.2 STL Baselines
Appendix H Details of RAA & RR Experiments: Hopper
Table 2: Hyperparameters for Hopper Learning
Hyperparameters for DOHJ-PPO Values
@@ Appendix H Details of RAA & RR Experiments: Hopper
Number of Environments 128
Number of Steps 400
Total Timesteps (RAA) 50M
Total Timesteps (RR) 50M
Scan Steps 4
Update Epochs 10
Number of Minibatches 32
@@
The Hopper environment is taken from Gym Brockman et al. [2016] and So et al. [2024]. In both RAA and RR problems, we define rewards and penalties based on the position of the Hopper head, which we denote as $(x,y)$ in this section.
In the RAA task, the reward is defined as
$r(x,y)=\sqrt{|x-2|+|y-1.4|}-0.1$ (54)
In the RR task, the first reward is defined again as
$r_{1}(x,y)=\sqrt{|x-2|+|y-1.4|}-0.1$ (55)
$r_{2}(x,y)=\sqrt{|x-0|+|y-1.4|}-0.1$ (56)
In all experiments, the Hopper is initialized in the default standing posture at a random $x\in[0,2]$ so as to learn a position-agnostic policy. The DOHJ-PPO parameters used to train these problems can be found in Table 2.
Appendix I Details of RAA & RR Experiments: F16
The F16 environment is taken from So et al. [2024], including a F16 fighter jet with a 26 dimensional observation. The jet is limited to a flight corridor with up to $2000$ relative position north ($x_{PN}$), $1200$ relative altitude ($x_{H}$), and $\pm 500$ relative position east ($x_{PE}$).
@@ Appendix I Details of RAA & RR Experiments: F16
In all experiments, the F16 is initialized with position $x_{PN}\in[250,750],x_{H}\in[300,900],x_{PE}\in[-250,250]$ and velocity in $v\in[200,450]$. Additionally, the roll, pitch, and yaw are initialized with $\pm\pi/16$ to simulate a variety of approaches to the flight corridor. Further details can be found in So et al. [2024]. The DOHJ-PPO parameters used to train these problems can be found in Table 3.
Table 3: Hyperparameters for F16 Learning
Hyperparameters for DOHJ-PPO Values
@@
Number of Environments 256
Number of Steps 200
Total Timesteps (RAA) 50M
Total Timesteps (RR) 100M
Scan Steps 10
Update Epochs 10
Number of Minibatches 64
@@
Appendix J Details of RAA & RR Experiments: SafetyGym
The SafetyGym environment is taken from the SafetyGym Point Environment Brockman et al. [2016], reimplemented in So et al. [2024]. In both RAA and RR problems, we define rewards and penalties based on the position of the Point position, which we denote as $(x,y)$ in this section.
In the RAA task, the reward is defined as
$r(x,y)=\sqrt{|x|+|y|}-0.3$ (60)
to incentive the Point to reach to the position at $(x,y)=(0.,0.)$. The penalty $q$ is defined as the minimum of signed distance functions to wall obstacles at $|x|\geq 3$ and $y\geq|3|$, and a scatter of eight random obstacles at $\{(1.4,0.6),(0.4,1.2),(-1.2,0.8),(-0.9,0.1.5),(-0.1,-1.1),(0.7,0.2),(-0.3,0.8),(-1.3,-1.3)\}$. In order to safely arrive at high reward (and always avoid the obstacles), the Point thus must drive through the obstacles and without leaving the box or crashing after arriving.
In the RR task, the first reward is defined again as
$r_{1}(x,y)=\sqrt{|x-2.5|+|y-2.5|}-0.3$ (61)
to incentive the Point to drive to the position at $(x,y)=(2.5,2.5)$, and the second reward as
$r_{2}(x,y)=\sqrt{|x+2.5|+|y+2.5|}-0.3$ (62)
to incentive the Point to drive to the position at (x,y)=(2.5,2.5)(x,y)=(-2.5,-2.5). In order to achieve both rewards, the Point must thus drive to both targets.
In all experiments, the Point is initialized at a random $(x,y)\in[-2,2]^{2}$ so as to learn a position-agnostic policy. The DOHJ-PPO parameters used to train these problems can be found in Table 4.
Table 4: Hyperparameters for SafetyGym Learning
Hyperparameters for DOHJ-PPO Values
Network Architecture MLP
Units per Hidden Layer 256
Numbers of Hidden Layers 2
Hidden Layer Activation Function tanh
Entropy coefficient Linear Decay 5e-3 $\rightarrow$ 0
Optimizer Adam
Discount factor $\gamma$ Linear Anneal 0.995 $\rightarrow$ 0.9995
GAE lambda parameter 0.95
Clip Ratio 0.2
Actor Learning rate Linear Decay 3e-4 $\rightarrow$ 0
Reward/Cost Critic Learning rate Linear Decay 3e-4 $\rightarrow$ 0
Number of Environments 128
Number of Steps 400
Total Timesteps (RAA) 50M
Total Timesteps (RR) 100M
Scan Steps 4
Update Epochs 10
Number of Minibatches 32
Add’l Hyperparameters for CPPO
$K_{P}$ 1
$K_{I}$ 1e-4
$K_{D}$ 1
Appendix K Details of RAA & RR Experiments: HalfCheetah
The HalfCheetah environment is taken from Gym Brockman et al. [2016], reimplemented in So et al. [2024]. In both RAA and RR problems, we define rewards based on the HalfCheetah head position, which we denote as $(x,y)$ in this section, and we define the penalties based on the position of all HalfCheetah body positions (back, neck, head, thighs, shins, and feet).
In the RAA task, the reward is defined as
r(x,y)=|x5.5|0.25r(x,y)=|x-5.5|-0.25 (63)
to incentive the HalfCheetah to reach its head to the cylinder centered at $x=5.5$. The penalty $q$ is defined as the minimum of signed distance functions to a floor obstacles at $(0,4.5)$ and $(0,6.5)$ which have height of $0.05$ and width $0.5$, and a back wall obstacle at x<0.9x<-0.9. In order to safely arrive at high reward (and always avoid the obstacles), the HalfCheetah thus must jump over the front floor obstacle (at $x=4.5$) and avoid crashing into, either by hopping over or standing, the back floor obstacle (at $x=6.5$).
In the RR task, the first reward is defined again as
$r_{1}(x,y)=\sqrt{|x-5|+|y-1|}-0.1$ (64)
to incentive the HalfCheetah to reach its head to the position at $(x,y)=(5,1)$, and the second reward as
$r_{2}(x,y)=\sqrt{|x+5|+|y-1|}-0.1$ (65)
to incentive the HalfCheetah to reach its head to the position at (x,y)=(5,1)(x,y)=(-5,1). In order to achieve both rewards, the HalfCheetah must thus gallop both forwards and backwards.
In all experiments, the HalfCheetah is initialized in the default standing posture at a random $x\in[0.5,1.5]$ so as to learn a position-agnostic policy. The DOHJ-PPO parameters used to train these problems can be found in Table 5.
Table 5: Hyperparameters for HalfCheetah Learning
Hyperparameters for DOHJ-PPO Values
Network Architecture MLP
Units per Hidden Layer 256
Numbers of Hidden Layers 2
Hidden Layer Activation Function tanh
Entropy coefficient Linear Decay 5e-3 $\rightarrow$ 0
Optimizer Adam
Discount factor $\gamma$ Linear Anneal 0.995 $\rightarrow$ 0.9995
GAE lambda parameter 0.95
Clip Ratio 0.2
Actor Learning rate Linear Decay 3e-4 $\rightarrow$ 0
Reward/Cost Critic Learning rate Linear Decay 3e-4 $\rightarrow$ 0
Number of Environments 128
Number of Steps 400
Total Timesteps (RAA) 100M
Total Timesteps (RR) 150M
Scan Steps 4
Update Epochs 10
Number of Minibatches 32
Add’l Hyperparameters for CPPO
$K_{P}$ 100
$K_{I}$ 1e-4
$K_{D}$ 1
Appendix L Contrast with Decompositions in Temporal Logic
Temporal Logic (TL) predicates, including those for reach (eventually $\varphi$), avoid (always not $\psi$), and reach-avoid (not $\psi$ until $\varphi$) problems, enjoy useful algebraic properties. These properties provide a convenient way to decompose complex TL predicates into simpler predicates. Unfortunately, the decomposition of these predicates does not generally translate to valid decompositions of the optimal value functions in HJR.
More explicitly, it is indeed possible to phrase HJR problem formulations, such as our RR and RAA problems, using the language of quantitative semantics from the TL literature. In particular, the standard HJR problem is to find the sequence of actions that maximizes some objective function, and this objective function can generally be written as a quantitative semantic corresponding to the specification of the desired system behavior (however, this is not explicitly written in TL notation in the HJR literature). That said, the optimal value function decomposition results for the RR and RAA problems are distinct from the decomposition of the corresponding quantitative semantics. In other words, our decomposition results are statements about the optimal control associated with the quantitative semantic, not the quantitative semantic itself.
This subtle distinction can be clarified by the following counter-examples, where a valid decomposition of the quantitative semantic (on the TL side) for the RR and RAA problem does not translate to a valid decomposition of the optimal value functions (on the HJR side) for the RAA problem. We will use $F$ as the eventually operator, $G$ as the always operator, and $\land$$\lor/$$\lnot$ as logical and/or/not.
L.1 RAA case
Consider an RAA problem where my friend is holding a piñata which I would like to break with a bat. We can always decompose the RAA quantitative semantic into R and A quantitative semantics. However, suppose there is some state of the system from which I can either eventually hit the piñata or always avoid hitting my friend, but not both. In this case, the optimal value function for the R and A problems will both be non-negative, even though I cannot actually achieve the RAA task from this state.
To make this argument explicit, define the atomic predicates $\varphi$ and $\psi$ to represent hitting the piñata and hitting my friend, respectively:
$(\mathbf{x},t)\models\varphi\iff r(\mathbf{x}(t))\geq 0,$
$(\mathbf{x},t)\models\psi\iff p(\mathbf{x}(t))\geq 0.$
Given a predicate $\mu$, let $\rho_{\mu}$ be the corresponding quantitative semantic. Thus, $\rho_{\varphi}[\mathbf{x},t]=r(\mathbf{x}(t))$ and $\rho_{\psi}[\mathbf{x},t]=p(\mathbf{x}(t))$. The quantitative semantics for the R, A, and RAA problems are then, respectively:
$\rho_{\mathrm{R}}[\mathbf{x},t]$ $:=\rho_{F\varphi}[\mathbf{x},t]=\max_{\tau\geq t}r(\mathbf{x}(\tau)),$ (66)
$\rho_{\mathrm{A}}[\mathbf{x},t]$ $:=\rho_{G\lnot\psi}[\mathbf{x},t]=\min_{\tau\geq t}-p(\mathbf{x}(\tau)),$ (67)
$\rho_{\mathrm{RAA}}[\mathbf{x},t]$ $:=\rho_{(F\varphi)\land(G\lnot\psi)}[\mathbf{x},t]=\min\left\{\max_{\tau\geq t}r(\mathbf{x}(\tau)),\min_{\tau\geq t}-p(\mathbf{x}(\tau))\right\}.$ (68)
As mentioned earlier, the optimal value functions for the R, A, and RAA problems can then be written in terms of the quantitative semantics, i.e.
$V_{\mathrm{R}}^{*}(s)$ $=\max_{\pi}\rho_{\mathrm{R}}[\mathbf{x}_{s}^{\pi},0],$
$V_{\mathrm{A}}^{*}(s)$ $=\max_{\pi}\rho_{\mathrm{A}}[\mathbf{x}_{s}^{\pi},0],$
$V_{\mathrm{RAA}}^{*}(s)$ $=\max_{\pi}\rho_{\mathrm{RAA}}[\mathbf{x}_{s}^{\pi},0],$
where $\mathbf{x}_{s}^{\pi}$ is the trajectory of the bat corresponding to the initial state $s$ and policy $\pi$.
But here is the key point: although it is always true that
$\rho_{\mathrm{RAA}}[\mathbf{x},t]=\min\{\rho_{\mathrm{R}}[\mathbf{x},t],\rho_{\mathrm{A}}[\mathbf{x},t]\},$
in general we only have that
$V_{\mathrm{RAA}}^{*}(s)\leq\min\{V_{\mathrm{R}}^{*}(s),V_{\mathrm{A}}^{*}(s)\}.$
Indeed this inequality is sometimes strict. Again, consider the case where there is some state of the system $s$ from which I can either eventually hit the piñata or always avoid hitting my friend, but cannot do both. Then
$V_{\mathrm{RAA}}^{*}(s)<0\leq\min\{V_{\mathrm{R}}^{*}(s),V_{\mathrm{A}}^{*}(s)\}.$
In other words, the algebra that applies to the TL predicates translates to an algebra on the TL quantitative semantics, but it does not generally translate to an algebra on the optimal value functions. This is why our decomposition required thorough justification.
L.2 RR case
Next, suppose that we are solving an RR problem for a small robotic boat that must deliver supplies to two downstream islands in a wide river that flows from north to south. The islands are at the same latitude, but one is to the west and one is to the east. Suppose the boat starts far enough upstream that it can reach either island, but the current is too strong to move from one island to the other. Then the boat can satisfy the R task for the west island (by ignoring the east island) and it can satisfy the R task for the east island (by ignoring the west island), but it cannot satisfy the RR task. The decomposition of the quantitative semantics for the RR problem into the minimum of the quantitative semantics for the two R problems will then not translate to a valid decomposition of the value functions.
More explicitly, with $r_{1}$ and $r_{2}$ the signed distance functions to either island, let
$(\mathbf{x},t)$ $\models\varphi\iff r_{1}(\mathbf{x}(t))\geq 0,$ (69)
$(\mathbf{x},t)$ $\models\psi\iff r_{2}(\mathbf{x}(t))\geq 0.$ (70)
Thus, $\rho_{\varphi}[\mathbf{x},t]=r_{1}(\mathbf{x}(t))$ and $\rho_{\psi}[\mathbf{x},t]=r_{2}(\mathbf{x}(t))$. The quantitative semantics for reaching the west island, reaching the east island, and reaching both islands are then respectively,
$\rho_{\mathrm{R1}}[\mathbf{x},t]$ $:=\rho_{F\varphi}[\mathbf{x},t]=\max_{\tau\geq t}r_{1}(\mathbf{x}(\tau)),$ (71)
$\rho_{\mathrm{R2}}[\mathbf{x},t]$ $:=\rho_{F\psi}[\mathbf{x},t]=\max_{\tau\geq t}r_{2}(\mathbf{x}(\tau)),$ (72)
$\rho_{\mathrm{RR}}[\mathbf{x},t]$ $:=\rho_{(F\varphi)\land(F\psi)}[\mathbf{x},t]=\min\left\{\max_{\tau\geq t}r_{1}(\mathbf{x}(\tau)),\max_{\tau\geq t}r_{2}(\mathbf{x}(\tau))\right\}.$ (73)
As before, the optimal value functions for the two R problems and the RR problems can then be written in terms of the quantitative semantics, i.e.
$V_{\mathrm{R1}}^{*}(s)$ $=\max_{\pi}\rho_{\mathrm{R1}}[\mathbf{x}_{s}^{\pi},0],$ (74)
$V_{\mathrm{R2}}^{*}(s)$ $=\max_{\pi}\rho_{\mathrm{R2}}[\mathbf{x}_{s}^{\pi},0],$ (75)
$V_{\mathrm{RR}}^{*}(s)$ $=\max_{\pi}\rho_{\mathrm{RR}}[\mathbf{x}_{s}^{\pi},0].$ (76)
But we still have the analogous issue: although it is always true that
$\rho_{\mathrm{RR}}[\mathbf{x},t]=\min\{\rho_{\mathrm{R1}}[\mathbf{x},t],\rho_{\mathrm{R2}}[\mathbf{x},t]\},$
in general we only have that
$V_{\mathrm{RR}}^{*}(s)\leq\min\{V_{\mathrm{R1}}^{\*}(s),V_{\mathrm{R2}}^{*}(s)\}.$
When we begin from some state of the system $s$ from which I can eventually reach the east island or I can eventually reach the west island, but not both, then the inequality is strict:
$V_{\mathrm{RR}}^{*}(s)<0\leq\min\{V_{\mathrm{R1}}^{\*}(s),V_{\mathrm{R2}}^{*}(s)\}.$
Appendix M Broader Impacts
@@
Appendix N Acknowledgments
We would like to thank several people for assisting the development of the work. We thank Jaime Fisac and Kai-Chieh Hsu for lending their knowledge on bridging Hamilton-Jacobi reachability and reinforcement learning. We thank Chuchu Fan and Oswin so for offering their knowledge on PPO-based HJ-RL and algorithm design. Lastly, we thank Sean Gao and Nikolay Atanasov for useful discussions regarding the paper.